# Towards autonomous photogrammetric forest inventory using a lightweight under-canopy robotic drone

### v2

@@ Abstract
Drones are increasingly used in forestry to capture high-resolution remote sensing data, supporting enhanced monitoring, assessment, and decision-making processes. While operations above the forest canopy are already highly automated, flying inside forests remains challenging, primarily relying on manual piloting. Inside dense forests, reliance on the Global Navigation Satellite System (GNSS) for localization is not feasible. Additionally, the drone must autonomously adjust its flight path to avoid collisions. Recently, advancements in robotics have enabled autonomous drone flights in GNSS-denied obstacle-rich areas. In this article, a step towards autonomous forest data collection is taken by building a prototype of a robotic under-canopy drone utilizing state-of-the-art open-source methods and validating its performance for data collection inside forests. Specifically, the study focused on camera-based autonomous flight under the forest canopy and photogrammetric post-processing of the data collected with the low-cost onboard stereo camera. The autonomous flight capability of the prototype was evaluated through multiple test flights at boreal forests. The tree parameter estimation capability was studied by performing diameter at breast height (DBH) estimation. The prototype successfully carried out flights in selected challenging forest environments, and the experiments showed excellent performance in forest 3D modeling with a miniaturized stereoscopic photogrammetric system. The DBH estimation achieved a root mean square error (RMSE) of 3.33 cm (12.79 %) across all trees. For trees with a DBH less than 30 cm, the RMSE was 1.16 cm (5.74 %). The results provide valuable insights into autonomous under-canopy forest mapping and highlight the critical next steps for advancing lightweight robotic drone systems for mapping complex forest environments.
@@
articletype: ARTICLE PREPRINT - Word Count 10048
@@ 1 Introduction
Over the last decades, remote sensing measurements within forests have been increasingly used to capture detailed forest information. Terrestrial laser scanning (TLS) made a breakthrough in the early 2000s, followed by the development of various solutions, including static laser scanning using tripods (Bienert et al. 2007) and mobile solutions, e.g., handheld devices (Ryding et al. 2015; Bauwens et al. 2016; Marselis et al. 2016; Balenović et al. 2021), backpacks (Liang et al. 2014b, 2018; Hyyppä et al. 2020b), ground-based vehicles (Forsman, Holmgren, and Olofsson 2016; Liang et al. 2018; Bienert et al. 2018) or drones (Chisholm et al. 2013; Brede et al. 2017; Wieser et al. 2017; Hyyppä et al. 2020a). Similarly, the use of structure-from-motion (SFM) based point clouds for forest mapping has been demonstrated using different camera solutions (e.g., Liang et al. 2014a; Wallace et al. 2016; Kuželka and Surovỳ 2018; Piermattei et al. 2019). Among these methods, drone-based solutions offer several benefits, including the ability to move in areas with challenging terrain and the high potential for automatization.
@@
Hence, the use of drones in forestry research has rapidly increased, and they offer a practical solution for forest data collection above the forest canopy (Puliti et al. 2015; Tuominen et al. 2017). Modern commercial drones are highly automated and can execute data collection missions highly autonomously in open-air environments using autopilot systems. However, their positioning systems typically rely exclusively on GNSS signals, limiting autonomous flights to open areas where these signals are reliable. In dense forests, GNSS localization becomes unreliable due to signal blockages and multipath effects caused by trees (Schubert et al. 2010). Additionally, navigating inside forests requires the capability to autonomously avoid obstacles such as trees and bushes along the flight path. Consequently, drone data collection flights under the forest canopy are still primarily controlled manually by human pilots (e.g., Hyyppä et al. 2020a; Liang et al. 2019; Krisanski, Taskhiri, and Turner 2020; Prabhu et al. 2024).
@@
Recent advances in robotics and the enhanced computational capabilities of small embedded computers have enabled drone autonomy even in GNSS-denied, cluttered environments. Although technology is still in the early stage of development, some recent studies have proposed various methods that have been successfully tested in forest environments (e.g., Liu et al. 2022; Loquercio et al. 2021; Zhou et al. 2022). However, the proposed methods are typically validated by reporting a single successful test flight in a specific forest environment. Most studies assess the effects of varying object density only in simulators, with only a couple of the methods tested through multiple test flights inside real forests.
@@
Moreover, most of the studies proposing methods for autonomous under-canopy drone trajectory planning have the scope solely on aerial robotics, and the sensor data is primarily utilized only for navigation. Only a few pioneering studies have yet explored the potential of custom robotic under-canopy drone systems in forestry applications (Liu et al. 2022; Liang et al. 2024; Cheng et al. 2024). Notably, all of these studies have utilized Ouster OS1 3D LiDAR (Ouster, San Francisco, CA, USA) as the main sensor. Whereas LiDARs have their strengths, such as 360° horizontal field of view (FOV), long sensing range, and robustness against varying illumination, they tend to be heavier and have higher power consumption than cameras (Reddy Cenkeramaddi et al. 2020). In the case of drone development, where optimizing the weight and battery consumption play a key role, these factors lead to bigger and heavier drone platforms and flight batteries requiring more space between the obstacles to fly safely inside the forest. Moreover, LiDARs with suitable quality are more expensive than camera systems with built-in 3D imaging capability, leading to increased building cost. The data collected with the LiDARs is also limited to point clouds only, whereas the images collected with cameras can also be used for image-based analysis.
@@
Hence, although some LiDAR-based robotic drone systems have been successfully applied to under-canopy mapping of forest parameters, the potential of smaller and cheaper camera-based systems remains largely unexplored. Furthermore, the robotic drone systems proposed in the literature are typically not evaluated through multiple flight tests in different forest environments leaving the reliability and suitability to diverse forest environments as an open question.
@@
Specifically, the study focused on three research questions. What level of autonomous flight reliability can be achieved in challenging boreal forests with current methods for camera-based robotic drone navigation? What level of DBH estimation accuracy can be achieved with the low-cost onboard stereo-camera data recorded during the flight? What are the remaining bottlenecks to be tackled in follow-up studies in order to achieve large-scale autonomous photogrammetric forest data collecting with miniaturized under-canopy robotic drone systems?
@@
  • Combining autonomous flight capabilities to under-canopy photogrammetric forest reconstruction by processing the data from a low-cost onboard stereo-camera to 3D point clouds.
@@
The rest of this study is structured as follows. Section 2 reviews the latest studies that estimate DBH using drones flying inside forests and recent open-source solutions for autonomous drone flights inside forests. Section 3 describes the algorithms, hardware, experiments, and methods utilized in this study. Experimental results are presented in Section 4. Section 5 discusses the results, and potential improvements, and concludes the study.
@@ 2.1 DBH estimation from under-canopy drone data
In a forest inventory, DBH is one of the most important attributes to estimate, as it provides a measure of the size of a tree and enables the prediction of the total stem volume either alone or combined with the three height (Laasasenaho 1982). Although some of the relevant studies have also estimated tree volume by flying the flight path multiple times with different altitudes, with rotating sensors, or with Lidars with high vertical FOV, only DBH estimation accuracies and methods are included in this brief literature review.
@@
In the existing literature, LiDAR-based methods have been studied extensively for the purpose of efficiently and accurately estimating various tree attributes, such as DBH, tree height, and stem volume. To date, several LiDAR-based mobile platforms have been shown to enable DBH estimation of individual trees with an RMSE of 1-2 cm (Ryding et al. 2015; Bauwens et al. 2016; Cabo et al. 2018; Chen et al. 2019; Hyyppä et al. 2020b, a, 2022), though most studies have focused on relatively sparse forests with little understory vegetation.
@@
Chisholm et al. (2013) were the first to demonstrate DBH measurements from point cloud data collected with a manually piloted drone flying under the forest canopy. The utilized custom-built drone system included Hokuyo UTM-30LX 2D LiDAR (Hokuyo, Osaka, Japan). They detected 12 trees in the vicinity of the drone trajectory and reported an RMSE of 25.1 % for the DBH estimates of the detected trees.
@@
Hyyppä et al. (2020a) collected high-quality point cloud data using a Kaarta Stencil-1 (Kaarta, Pittsburgh, Pennsylvania, USA) laser scanner mounted horizontally at the bottom of a manually piloted drone. They demonstrated DBH estimates with an RMSE of 0.69-0.92 cm (2.2 %-3.1 %) and a bias of 0.29-0.31 cm (1.0 %-1.1 %) on two plots of varying complexity in a boreal forest with a tree density of 410-420 trees/ha. Importantly, their stem diameter estimation algorithm utilized the temporal order of data points to detect stem arcs, thus mitigating distortions caused by slow temporal drifts of the drone trajectory. In a follow-up study, Hyyppä et al. (2020c) compared various mobile laser-scanning methods, including backpack, handheld, Under-Canopy UAV, and Above-Canopy UAV, for forest data collection. The same manually piloted under-canopy drone system was utilized in this study in the same forest plots as in Hyyppä et al. (2020a). In this study, the manually piloted under-canopy drone system yielded an RMSE of 0.6-1.1 cm (2.3 %-3.5 %) and a bias of 0.34-0.12 cm (1.3 %-0.4 %). Wang et al. (2021) combined data from above- and under-canopy flights of a manually-piloted drone, resulting in an RMSE of 7.95 cm (33.35 %) and a bias of 3.72 cm (15.61 %) for the DBH.
@@
Muhojoki et al. (2024) compared the accuracy of the estimated tree attributes from several LiDAR-based mobile systems, including two under-canopy drone systems, a system by Hovermap (2020 version, Emesent, Milton, Australia) equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA) and a system by Deep Forestry (Deep Forestry, Uppsala, Sweden) utilizing an Ouster OS0-32 Rev. 5 laser scanner (Ouster, San Francisco, CA, USA). Deep Forestry drone was flown manually, but the Hovermap drone system was able to fly autonomously in sparse forests where the distance between the trees was at least 3 m. Both drone systems were used to collect point clouds under the canopy of three sparse forest plots. The RMSE and bias of DBH estimates for Deep Forestry and Hovermap drones, with bias removal method introduced by Muhojoki et al. (2024), averaged RMSE of 1.5 cm (6.7 %) and bias of -0.7 cm (3.0 %) and RMSE of 1.6 cm (6.9 %) and bias of -0.1 cm (-0.6 %), respectively. Without using the bias removal method, RMSE and bias for both drones were significantly higher.
@@
In addition to the manually controlled drones and large autonomous commercial drone systems, recently a few pioneering studies have applied smaller custom robotic drone platforms with a high automation level for LiDAR-based under-canopy DBH estimation. Liang et al. (2024) applied an autonomously operated under-canopy drone system to collect laser scanning data inside a 50 m $\times$ 30 m forest plot with a tree density of approximately 1000 trees/ha. To enable autonomous flight, a predefined global path was planned, to which the drone system performed real-time adjustments during the flight according to real-time environmental conditions. When the flight path covered most of the test plot, they achieved an RMSE of 5.13 cm (22.01 %) and bias of 3.40 cm (14.62 %) for the DBH of individual trees.
@@
Liu et al. (2022) employed a LiDAR-based semantic lidar odometry and mapping framework (SLOAM) (Chen et al. 2020) algorithm to calculate the number of trees in the explored area. Although in this study the autonomous flight was demonstrated without the SLOAM algorithm and the tree number calculation was only demonstrated with one manually-piloted test flight, the same drone system has later been used in various follow-up studies for tree parameter estimation. Prabhu et al. (2024) extended the system to have 3D trajectory planning capability and used the platform to demonstrate metric-semantic map merging between multiple robotic drones as well as DBH and tree trunk diameter profile estimation with the onboard LiDAR data. The DBH values were estimated for 20 trees located in two different test plots with RMSE of 1.67 cm (7.42 %). However, the DBH data-collecting flights were performed with manual flight mode. Cheng et al. (2024) used the same drone platform to collect a dataset for LiDAR-based mapping of trees. In addition to data from orchards, the dataset contains the onboard LiDAR data from two forest areas. Although the drone was manually piloted in one of these forests, the data collection at the other forest consisting of loblolly pine trees was performed in autonomous flight mode. They also estimated the DBH value for 77 trees from the LiDAR data collected during the autonomous flights achieving an RMSE of 3.8 cm (12.3 %) by offline afterprocessing and 9.8 cm (31.8 %) by online estimation with SLOAM. With manually piloted data collection in the second test forest consisting of 20 trees with varying species, the RMSEs were 2.0 cm (7.5 %) with offline processing and 7.5 cm (27.5 %) with online processing.
Besides the under-canopy flying drones, LiDARs have also been integrated to above-canopy flying drones capable of direct stem diameter measurements, though typically with a significantly higher RMSE and a lower tree detection rate due to the larger measurement distance and the upper canopy obstructing the tree stems (Wieser et al. 2017; Brede et al. 2017; Kuželka, Slavík, and Surovỳ 2020; Vandendaele et al. 2021; Feng et al. 2022).
@@
During the past decade, photogrammetric systems producing SfM point clouds have been suggested to provide an efficient, low-cost alternative to traditional field measurements of tree attributes, such as DBH. Even though several studies have investigated terrestrial photogrammetric systems (e.g., Liang et al. (2014a); Surovỳ, Yoshimoto, and Panagiotidis (2016); Mokroš et al. (2018); Piermattei et al. (2019); Xu, Shen, and Cao (2023); Gao and Kan (2022)) and above-canopy photogrammetric systems (e.g., Wallace et al. (2016); Ye, van Leeuwen, and Nyktas (2019)) for tree attribute estimation, only relatively few studies have tested the use of under-canopy drone systems for the collection of SfM point clouds enabling tree attribute measurements (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; Shimabuku, Konoshima, and Ota 2023; He et al. 2025).
@@
Kuželka and Surovỳ (2018) manually operated a drone system based on a DJI Phantom 4 Pro (DJI, Shenzhen, China) within two forest plots to produce low-noise photogrammetric point clouds, from which the DBH of individual trees could be estimated with an RMSE between 2.6 cm and 5.3 cm depending on the plot and the applied analysis method. The studied plots were sparse with only 270-290 trees/ha. Krisanski, Taskhiri, and Turner (2020) performed manual flights with a DJI Phantom 4 Pro inside two forest plots with a stem density of approximately 1000 trees/ha to acquire high-density SfM point cloud data, from which the DBH of individual trees could be measured with an RMSE of 1.1-2.1 cm.
@@
Shimabuku, Konoshima, and Ota (2023) achieved an RMSE of 0.4-0.7 cm for DBH estimation on two small subtropical forest plots using SfM point clouds acquired with an under-canopy flying DJI Mini 3 Pro drone (DJI, Shenzhen, China). The first plot containing 18 trees was clean containing a minimal amount of low-hanging branches and understory vegetation. The second test plot was denser and closer to a natural subtropical forest environment but contained only 7 trees. The drone system was able to detect obstacles in the horizontal direction but needed manual piloting to avoid obstacles in the vertical direction and to guide the drone back to the intended flight path after automatic obstacle avoidance in the horizontal direction.
@@
Recently, He et al. (2025) mounted an additional 360° FOV Insta360 camera (Arashi Vision, Shenzhen, China) to the same LiDAR-based autonomous drone platform used in Liu et al. (2022) and post-processed the photogrammetric data to 3D point clouds. Whether the flights were performed in autonomous flight mode or with manually piloted mode was not mentioned in the study. Their study evaluated DBH for 43 trees with varying tree species and the reported median relative DBH error was 11.5 % estimated from the photogrammetric point clouds.
A summary of the results obtained in the relevant studies on under-canopy forest inventory is presented in Table 1. In summary of the related work, most of the studies for under-canopy drone-based DBH estimation have been performed with manually piloted drones. Although a few studies have explored the potential of custom robotic drone systems capable of flying without commands from the pilot during the flight, all of these systems have been based on LiDARs. All of the studies based on photogrammetry and SfM point clouds have utilized commercial drone systems, which have required manual piloting during the data collection.
@@
The highest estimation accuracies have been obtained with large systems unable to fly in dense forests or with test plots containing very few reference trees, but miniaturized systems have still yielded promising results. Notably, all studies utilizing SfM point clouds have been performed in environments that differ significantly from boreal forests.

### v3

@@ Abstract
Drones are increasingly used in forestry to capture high-resolution remote sensing data, supporting enhanced monitoring, assessment, and decision-making processes. While operations above the forest canopy are already highly automated, flying inside forests remains challenging, primarily relying on manual piloting. In dense forests, relying on the Global Navigation Satellite System (GNSS) for localization is not feasible. In addition, the drone must autonomously adjust its flight path to avoid collisions. Recently, advancements in robotics have enabled autonomous drone flights in GNSS-denied obstacle-rich areas. In this article, a step towards autonomous forest data collection is taken by building a prototype of a robotic under-canopy drone utilizing state-of-the-art open source methods and validating its performance for data collection inside forests. Specifically, the study focused on camera-based autonomous flight under the forest canopy and photogrammetric post-processing of the data collected with the low-cost onboard stereo camera. The autonomous flight capability of the prototype was evaluated through multiple test flights in boreal forests. The tree parameter estimation capability was studied by performing diameter at breast height (DBH) estimation. The prototype successfully carried out flights in selected challenging forest environments, and the experiments showed promising performance in forest 3D modeling with a miniaturized stereoscopic photogrammetric system. The DBH estimation achieved a root mean square error (RMSE) of 3.33 - 3.97 cm (10.69 - 12.98 %) across all trees. For trees with a DBH less than 30 cm, the RMSE was 1.16 - 2.56 cm (5.74 - 12.47 %). The results provide valuable insights into autonomous under-canopy forest mapping and highlight the critical next steps for advancing lightweight robotic drone systems for mapping complex forest environments.
@@
articletype: ARTICLE PREPRINT
@@ 1 Introduction
In recent decades, remote sensing measurements within forests have been increasingly used to capture detailed forest information. Terrestrial laser scanning (TLS) made a breakthrough in the early 2000s, followed by the development of various solutions, including static laser scanning using tripods (Bienert et al. 2007) and mobile solutions, e.g., handheld devices (Ryding et al. 2015; Bauwens et al. 2016; Marselis et al. 2016; Balenović et al. 2021), backpacks (Liang et al. 2014b, 2018; Hyyppä et al. 2020b), ground-based vehicles (Forsman, Holmgren, and Olofsson 2016; Liang et al. 2018; Bienert et al. 2018) or drones (Chisholm et al. 2013; Brede et al. 2017; Wieser et al. 2017; Hyyppä et al. 2020a). Similarly, the use of structure-from-motion (SfM) based point clouds for forest mapping has been demonstrated using different camera solutions (e.g., Liang et al. 2014a; Wallace et al. 2016; Kuželka and Surovỳ 2018; Piermattei et al. 2019). Among these methods, drone-based solutions offer several benefits, including the ability to move in areas with challenging terrain and the high potential for automatization.
@@
Hence, the use of drones in forest research has increased rapidly and offers a practical solution to collect forest data above the forest canopy (Puliti et al. 2015; Tuominen et al. 2017). Modern commercial drones are highly automated and can execute data collection missions highly autonomously in open-air environments using autopilot systems. However, their positioning systems typically rely exclusively on GNSS signals, limiting autonomous flights to open areas where these signals are reliable. In dense forests, GNSS-based localization becomes unreliable due to signal blockages and multipath effects caused by trees (Schubert et al. 2010). Furthermore, navigating within forests requires the ability to autonomously avoid obstacles such as trees and bushes along the flight path. Consequently, drone data collection flights under the forest canopy are still primarily controlled manually by human pilots (e.g., Hyyppä et al. 2020a; Liang et al. 2019; Krisanski, Taskhiri, and Turner 2020; Prabhu et al. 2024).
@@
Recent advances in robotics and the enhanced computational capabilities of small embedded computers have enabled drone autonomy even in cluttered GNSS-denied environments. Although technology is still in the early stage of development, some recent studies have proposed various methods that have been successfully tested in forest environments (e.g., Liu et al. 2022; Loquercio et al. 2021; Zhou et al. 2022). However, the proposed methods are typically validated by reporting a single successful test flight in a specific forest environment. Most studies assess the effects of varying object density only in simulators, with only a couple of methods tested through multiple test flights inside real forests.
@@
Moreover, most of the studies proposing methods for autonomous under-canopy drone trajectory planning have the scope solely on aerial robotics, and the sensor data is primarily utilized only for navigation. Only a few pioneering studies have yet explored the potential of custom robotic under-canopy drone systems in forestry applications (Liu et al. 2022; Liang et al. 2024; Cheng et al. 2024). Notably, all of these studies have utilized Ouster OS1 3D LiDAR (Ouster, San Francisco, CA, USA) as the main sensor. Whereas LiDARs have their strengths, such as 360° horizontal field of view (FOV), long sensing range, and robustness against varying illumination, they tend to be heavier and have higher power consumption than cameras (Reddy Cenkeramaddi et al. 2020). In the case of drone development, where optimizing the weight and battery consumption plays a key role, these factors lead to larger and heavier drone platforms and flight batteries that require more space between obstacles to fly safely inside the forest. Furthermore, LiDARs with suitable quality are more expensive than camera systems with built-in 3D imaging capability, leading to increased building cost. The data collected with the LiDARs is also limited to point clouds only, whereas the images collected with cameras can also be used for image-based analysis.
@@
Hence, although some LiDAR-based robotic drone systems have been successfully applied to under-canopy mapping of forest parameters, the potential of smaller and cheaper camera-based systems remains largely unexplored. Furthermore, the robotic drone systems proposed in the literature are typically not evaluated through multiple flight tests in different forest environments, leaving the reliability and suitability to diverse forest environments as an open question.
@@
Specifically, the study focused on three research questions. What level of autonomous flight reliability can be achieved in challenging boreal forests with current methods for camera-based robotic drone navigation? What level of DBH estimation accuracy can be achieved with the low-cost onboard stereo camera data recorded during the flight? What are the remaining bottlenecks to be tackled in follow-up studies in order to achieve large-scale autonomous photogrammetric forest data collecting with miniaturized under-canopy robotic drone systems?
@@
  • Combining autonomous flight capabilities to under-canopy photogrammetric forest reconstruction by processing the data from a low-cost onboard stereo camera to 3D point clouds.
@@
The rest of this study is structured as follows. Section 2 reviews the latest studies that estimate DBH using drones flying inside forests and recent open source solutions for autonomous drone flights inside forests. Section 3 describes the algorithms, hardware, experiments, and methods utilized in this study. Experimental results are presented in Section 4. Section 5 discusses the results and potential improvements and concludes the study.
@@ 2.1 DBH estimation from under-canopy drone data
In a forest inventory, DBH is one of the most important attributes to estimate, as it provides a measure of the size of a tree and enables the prediction of the total volume of the stem alone or in combination with the three heights (Laasasenaho 1982).
@@
In the existing literature, LiDAR-based methods have been extensively studied to efficiently and accurately estimate various tree attributes, such as DBH, tree height, and stem volume. To date, several LiDAR-based mobile platforms have been shown to enable DBH estimation of individual trees with an RMSE of 1-2 cm (Ryding et al. 2015; Bauwens et al. 2016; Cabo et al. 2018; Chen et al. 2019; Hyyppä et al. 2020b, a, 2022), though most studies have focused on relatively sparse forests with little understory vegetation.
@@
Chisholm et al. (2013) were the first to demonstrate DBH measurements from point cloud data collected with a manually piloted drone flying under the forest canopy. The utilized custom-built drone system included a Hokuyo UTM-30LX 2D LiDAR (Hokuyo, Osaka, Japan).
@@
Hyyppä et al. (2020a) collected high-quality point cloud data using a Kaarta Stencil-1 (Kaarta, Pittsburgh, Pennsylvania, USA) laser scanner mounted horizontally at the bottom of a manually piloted drone in two plots of varying complexity in a boreal forest with a tree density of 410-420 trees/ha. Importantly, their stem diameter estimation algorithm utilized the temporal order of data points to detect stem arcs, thus mitigating distortions caused by slow temporal drifts of the drone trajectory. In a follow-up study, Hyyppä et al. (2020c) compared various mobile laser scanning methods, including backpack, handheld, Under-Canopy drone, and Above-Canopy drone, for forest data collection. In this study, the same manually piloted under-canopy drone system was utilized in the same forest plots as in Hyyppä et al. (2020a).
Wang et al. (2021) combined data from above and under canopy flights of a manually piloted drone equipped with a Riegl miniVUX laser scanner (RIEGL Laser Measurement Systems GmbH, Horn, Austria). The study area was a boreal forest plot with an approximate tree density of 200 trees/ha.
@@
Muhojoki et al. (2024) compared the accuracy of the estimated tree attributes of several LiDAR-based mobile systems, including two commercial under-canopy drone systems, a Hovermap system (2020 version, Emesent, Milton, Australia) equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA) and a Deep Forestry system (Deep Forestry, Uppsala, Sweden) utilizing an Ouster OS0-32 Rev. 5 laser scanner (Ouster, San Francisco, CA, USA). The Deep Forestry drone was flown manually, but the Hovermap drone system was able to fly autonomously in sparse forests where the distance between trees was at least 3 m. Both drone systems were used to collect point clouds under the canopy of three sparse forest plots.
@@
In addition to manually controlled drones and large autonomous commercial drone systems, recently a few pioneering studies have applied smaller custom robotic drone platforms with a high automation level for LiDAR-based under-canopy DBH estimation. Liang et al. (2024) applied an autonomously operated under-canopy drone system to collect laser scanning data inside a 50 m $\times$ 30 m forest plot with a tree density of approximately 1000 trees/ha. To enable autonomous flight, a predefined global path was planned, in which the drone system made real-time adjustments during the flight according to real-time environmental conditions.
@@
Liu et al. (2022) employed a LiDAR-based semantic lidar odometry and mapping framework (SLOAM) (Chen et al. 2020) algorithm to calculate the number of trees in the explored area. Although in this study autonomous flight was demonstrated without the SLOAM algorithm and tree number calculation was only demonstrated with one manually piloted test flight, the same drone system has later been used in various follow-up studies for tree parameter estimation. Prabhu et al. (2024) extended the system to have 3D trajectory planning capability and used the platform to demonstrate metric-semantic map merging between multiple robotic drones, as well as DBH and tree trunk diameter profile estimation with the onboard LiDAR data. However, DBH data collection flights were performed in manual flight mode. Cheng et al. (2024) used the same drone platform to collect a dataset for tree mapping based on LiDAR. In addition to data from orchards, the dataset contains the onboard LiDAR data from two forest areas. Although the drone was manually piloted in one of these forests, data collection in the other forest consisting of loblolly pine trees was carried out in autonomous flight mode. They also estimated DBH values for the LiDAR data collected during both autonomous and manual flights.
@@
During the past decade, photogrammetric systems producing SfM point clouds have been suggested to provide an efficient, low-cost alternative to traditional field measurements of tree attributes, such as DBH. Although several studies have investigated terrestrial photogrammetric systems (e.g., Liang et al. (2014a); Surovỳ, Yoshimoto, and Panagiotidis (2016); Mokroš et al. (2018); Piermattei et al. (2019); Xu, Shen, and Cao (2023); Gao and Kan (2022)) and above-canopy photogrammetric systems (e.g., Wallace et al. (2016); Ye, van Leeuwen, and Nyktas (2019)) for tree attribute estimation, only relatively few studies have tested the use of under-canopy drone systems for the collection of SfM point clouds enabling tree attribute measurements (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; Shimabuku, Konoshima, and Ota 2023; He et al. 2025).
Kuželka and Surovỳ (2018) evaluated two manually operated drone systems, a DJI Phantom 4 Pro (DJI, Shenzhen, China) and a DJI Mavic Pro (DJI, Shenzhen, China), and various analysis methods to estimate DBH from photogrammetric point clouds. The systems were evaluated in two sparse forest plots with tree densities of 270-290 trees/ha.
@@
Krisanski, Taskhiri, and Turner (2020) performed manual flights with a DJI Phantom 4 Pro in two forest plots with stem densities of 885 trees/ha and 980 to acquire high-density SfM point cloud data and estimated the DBH of individual trees from the point clouds.
@@
Shimabuku, Konoshima, and Ota (2023) performed DBH estimation on two small subtropical forest plots using SfM point clouds acquired with an under-canopy flying DJI Mini 3 Pro drone (DJI, Shenzhen, China). The drone system was able to detect obstacles in the horizontal direction, but needed manual piloting to avoid obstacles in the vertical direction and to guide the drone back to the intended flight path after automatic obstacle avoidance in the horizontal direction.
@@
Recently, He et al. (2025) mounted a 360° FOV Insta360 camera (Arashi Vision, Shenzhen, China) as an additional payload to the same LiDAR-based autonomous drone platform used in Liu et al. (2022) and post-processed the photogrammetric data to 3D point clouds. However, in the study, it was not mentioned whether the flights were performed in autonomous flight mode or with manual pilot mode.
The RMSEs of the DBH estimates in the relevant studies are presented in Table 1. From the studies where multiple analysis methods were evaluated, the lowest RMSEs are presented. The study by He et al. (2025) was omitted from the table, as the study reported only the median relative error (11.5 %) of the accuracy of the DBH estimation instead of the RMSE. Although some of the relevant studies also estimated tree volumes by flying the flight path several times with different altitudes, with rotating sensors, or with LiDARs with high vertical FOV, only DBH estimation accuracies are included in the table.
@@
In summary of the related work, most of the studies for under-canopy drone-based DBH estimation have been conducted with manually piloted drones. Although a few studies have explored the potential of custom robotic drone systems capable of flying without commands from the pilot during the flight, all of these systems have been based on LiDARs. All of the studies based on photogrammetry and SfM point clouds have utilized commercial drone systems, which have required manual piloting during data collection.
The highest estimation accuracies have been obtained with large systems unable to fly in dense forests or with test plots containing very few reference trees, but miniaturized systems have still yielded promising results. Furthermore, all studies utilizing SfM point clouds have been performed in environments that differ significantly from boreal forests.